# Clutter-Aware Target Detection for ISAC in a Millimeter-Wave Cell-Free Massive MIMO System

### v2

@@ Abstract
In this paper, we investigate the performance of an integrated sensing and communication (ISAC) system within a cell-free massive multiple-input multiple-output (MIMO) system. Each access point (AP) operates in the millimeter-wave (mmWave) frequency band. The APs jointly serve the user equipments (UEs) in the downlink while simultaneously detecting a target through dedicated sensing beams, which are directed toward a reconfigurable intelligent surface (RIS). Although the AP-RIS, RIS-target, and AP-target channels have both line-of-sight (LoS) and non-line-of-sight (NLoS) parts, it is assumed only knowledge of the LoS paths is available. A key contribution of this study is the consideration of clutter, which degrades the target detection if not handled. We propose an algorithm to alternatively optimize the transmit power allocation and the RIS phase-shift matrix, maximizing the target signal-to-clutter-plus-noise ratio (SCNR) while ensuring a minimum signal-to-interference-plus-noise ratio (SINR) for the UEs. Numerical results demonstrate that exploiting clutter subspace significantly enhances detection probability, particularly at high clutter-to-noise ratios, and reveal that an increased number of transmit side clusters impairs detection performance. Finally, we highlight the performance gains achieved using a dedicated sensing stream.
@@ I Introduction
Cell-free massive multiple-input multiple-output (MIMO) is a user-centric network infrastructure where a set of distributed access points serve multiple user equipments on the same time-frequency resources using joint processing techniques [1]. Over nearly a decade of research, it has been shown that cell-free massive MIMO improves both spectral and energy efficiency by providing macro diversity and enhanced interference management compared to traditional cellular setups. As a result, it has become one of the key technologies for sixth-generation (6G) wireless networks [2].
Recently, cell-free massive MIMO architecture has also demonstrated potential benefits in integrated sensing and communications (ISAC) [3, 4]. ISAC allows for the efficient use of hardware and spectral resources through integration and coordination gains, compared to separate communication and sensing architectures [5]. Cell-free massive MIMO facilitates multi-static sensing through a central processing unit (CPU), to which the APs are connected via fronthaul links. Beyond cell-free massive MIMO and ISAC, another key 6G technology is reconfigurable intelligent surface (RIS)-aided communications. RISs have attracted significant attention from both academia and industry due to their ability to control challenging radio propagation environments intelligently. By using low-cost passive reflecting elements that induce controllable phase shifts in incoming waveforms, RISs can steer reflected signals toward desired locations while reducing interference at undesired points [6, 7]. Furthermore, RISs have emerged as a promising technology in ISAC applications, enhancing the available spatial degrees-of-freedom. In [8], the sum rate of communication UEs is maximized under a worst-case sensing signal-to-noise ratio (SNR) constraint. An often overlooked aspect in ISAC research is the effect of clutter on target detection, which can severely degrade system performance if not properly addressed. Recent studies, such as [9], illustrate the impact of clutter in monostatic ISAC settings, underscoring the significance of clutter awareness at the receiving APs for reliable target detection.
@@
In this paper, we investigate the extent to which an RIS can enhance sensing capabilities when the direct channels between the transmitting APs and the target are obstructed. We also explore the need for precise channel state information (CSI) for target detection within a millimeter-wave (mmWave) channel architecture, as well as the requirement for sensing-specific streams. Our setup, illustrated in Fig. 1, evaluates the system’s ability to detect a potential target at a known location while meeting a minimum signal-to-interference-plus-noise ratio (SINR) for each UE. To this extent, we maximize the target’s signal-to-clutter-and-noise ratio (SCNR) using an alternating optimization (AO) algorithm that jointly optimizes transmit power allocation and the RIS phase-shift matrix. The APs assume a line-of-sight (LoS) structure for the sensing links, justified by the high path loss in mmWave frequencies, which typically makes LoS paths dominant over non-line-of-sight (NLoS) paths. A centralized precoding strategy is employed with regularized zero forcing (RZF) for communication streams. The sensing stream is projected onto the null space of the communication channels, which are assumed to be known at the APs. Assuming knowledge of the clutter subspace, target detection is conducted using a generalized likelihood ratio test (GLRT). The numerical simulations reveal that clutter awareness significantly enhances the detection probability, particularly at high clutter-to-noise ratio (CNR). Furthermore, when sensing links contain many clusters concentrated in a narrow angular range, the LoS assumption becomes less effective, reducing detection probability. The results also highlight the importance of allocating dedicated sensing symbols to improve detection performance.
Notation: $\odot$ represents the Hadamard (element-wise) product. Boldface lowercase and uppercase letters denote vectors and matrices. The trace of the matrix $\mathbf{X}$ is denoted by $\text{Tr}(\mathbf{X})$. $\text{diag}(\mathbf{x})$ represents the stacking of $\mathbf{x}$ on the main diagonal of a matrix. $\mathbf{I}_{N}$ represents an $N\times N$ identity matrix. The notation $\mathcal{CN}(0,\sigma^{2})$ represents circularly symmetric complex Gaussian distribution with variance $\sigma^{2}$. By $\mathcal{L}\setminus\{l\}$, we denote the set obtained by removing the element $l$ from the set $\mathcal{L}$. Given a matrix $\mathbf{M}$, $\mathbf{M}^{\dagger}$ represents its Moore-Penrose pseudoinverse.
@@ II System Model
We consider the RIS-assisted cell-free massive MIMO ISAC network shown in Fig. 1. Here a set $\mathcal{T}$ made of $T$ APs, each equipped with $M$ transmit antennas, serves $K$ single-antenna UEs in the downlink while sensing the potential presence of a target at a known position. The sensing is assumed to be carried out over $\tau$ symbols belonging to the same channel coherence block. A second set of $R$ APs, denoted by $\mathcal{R}$, collects the echos scattered from the target and performs target detection. The choice of geographically separated transmitting and receiving APs is motivated by the implementation challenges of full duplex operation, especially in terms of self interference [10], and the interest in this setup for cell-free massive MIMO communications. We assume that the direct channel between the APs in $\mathcal{T}$ and the target is blocked and, thus, the formers reach the target through a RIS equipped with $N$ reflective elements whose phase shifts are described by the vector $\boldsymbol{\theta}=[\theta_{1},\dots,\theta_{N}]^{\top}$, where $|\theta_{n}|=1$ for $n=1,\dots,N$. Each AP $t\in\mathcal{T}$ is connected to the RIS via the channel $\mathbf{H}_{t}\in\mathbb{C}^{N\times M}$, the reflected path connecting the RIS to the target is denoted by $\mathbf{c}\in\mathbb{C}^{N}$ and the UEs are connected to the APs in $\mathcal{T}$ via the channel $\{\mathbf{g}_{k,t}\}_{\forall k,t}\in\mathbb{C}^{M}$. The receiving APs in $\mathcal{R}$ are connected to the target via the channels $\{\mathbf{b}_{r}\}_{\forall r\in\mathcal{R}}\in\mathbb{C}^{M}$.
@@ II-A Channel models
Each of the previously mentioned channels is modeled in accordance to [11], which accurately captures the spatial sparsity that characterizes mmWave channels. Each channel consists of contributions from $C_{i}$ scattering clusters, i.e.,
@@
where the overline superscript indicates the LoS paths and the other terms correspond to the NLoS parts. The $\beta_{0,r}^{2}$, $\beta_{0}^{2}$, and $\beta_{0,t}^{2}$ are the corresponding LoS channel gains. We now remove the additional subscripts and superscripts except for $n$ as the following consideration applies regardless of the additional subscripts. Here, $\alpha_{n}$ is the complex channel gain of the $n$-th cluster. We assume each cluster scatters a sufficient number of rays such that $\alpha_{n}\sim\mathcal{CN}(0,\beta_{n}^{2})$ and the fading is mutually independent. The coefficients $\beta_{n}^{2}$ represent the path loss associated with the respective clusters. The geometry of the $n$-th cluster is described by $\boldsymbol{\omega}_{n}=[\psi_{n},\phi_{n}]$, where $\psi_{n}$ is the azimuth angle and $\phi_{n}$ is the elevation angle. Each AP is equipped with a uniform planar square array with a half-wavelength vertical and horizontal inter-element spacing such that the entries of $M$-dimensional steering vector are defined as
$[\mathbf{a}_{M}(\boldsymbol{\omega}_{n})]_{m_{h}m_{v}}=e^{-j\pi(m_{h}\sin(\psi_{n})\cos(\phi_{n})+m_{v}\sin(\phi_{n}))}$ (5)
for $m_{h},m_{v}=0,\dots,\sqrt{M}-1$. We assume that perfect CSI for $\{\mathbf{g}_{k,t}\}_{\forall k,t}$ is available at the APs, as this is easily obtainable through well-known estimation techniques. As for the sensing channels, the transmit APs only know the position of the RIS and the target and, thus, the corresponding LoS channels.
@@ II-B Communication observation model
with $\mathbf{s}_{t}[\iota]=\mathbf{F}_{t}\mathbf{X}[\iota]\boldsymbol{\rho}$, where the matrix $\mathbf{X}[\iota]=\text{diag}(x_{1}[\iota],\ldots,x_{L}[\iota])$ is a diagonal matrix containing the downlink communication symbols and sensing symbol transmitted during timeslot $\iota$, where $\mathbb{E}\{|x_{k}[\iota]|^{2}\}=1$. The vector $\boldsymbol{\rho}=[\sqrt{\rho_{1}},\dots,\sqrt{\rho_{L}}]^{\top}$ contains the square root of the power allocated to each beam. Finally, $w_{k}[\iota]\sim\mathcal{CN}(0,\sigma_{k}^{2})$. In line with the centralized operation of cell-free massive MIMO, the communication precoding vectors are based on the transmit APs’ CSI [3] while the sensing precoder is based on the assumed LoS structure of the sensing channels and the known target location. To this end, we define $\mathbf{g}_{k}=[\mathbf{g}_{k,1}^{\top},\dots,\mathbf{g}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$ and $\mathbf{f}_{k}=[\mathbf{f}_{k,1}^{\top},\dots,\mathbf{f}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$. The communication precoding vectors for each UE are chosen using RZF as
$\mathbf{f}_{k}=\frac{\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}}{\left\lVert\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}\right\rVert},\leavevmode\nobreak\ k=1,\dots,K,$ (7)
where $\lambda$ is a regularization parameter and this parameter allows the system to implement a trade-off between interference and noise suppression. For the sensing stream, let us define the matrix $\mathbf{G}=[\mathbf{g}_{1},\dots,\mathbf{g}_{K}]\in\mathbb{C}^{MT\times K}$ with all the UE channels. Then the sensing stream precoder is defined as
$\mathbf{f}_{L,t}=\frac{\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}}{\left\lVert\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}\right\rVert},$ (8)
where $\overline{\mathbf{h}}=[\overline{\mathbf{h}}_{1}^{\top},\dots,\overline{\mathbf{h}}_{T}^{\top}]^{\top}$ and $\overline{\mathbf{h}}_{t}=\overline{\mathbf{H}}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}$ with $\boldsymbol{\Theta}=\text{diag}(\boldsymbol{\theta})$. We have projected $\overline{\mathbf{h}}$ onto the null space of the communication UEs’ channels to null the interference caused to the UEs. The communication performance is represented by the UEs’ spectral efficiency (SE), defined as ${\rm{SE}}_{k}=\log_{2}(1+{\rm{SINR}}_{k})$: under the assumption of perfect CSI on the UEs channels, the SINR is defined as
@@ II-C Sensing observation model
$\mathbf{y}_{r}[\iota]=\sum_{t\in\mathcal{T}}\xi_{t,r}\mathbf{E}_{t,r}\mathbf{s}_{t}[\iota]+\mathbf{z}_{r}[\iota]+\mathbf{n}_{r}[\iota],$ (10)
where $\xi_{t,r}$ is the radar cross section (RCS) associated with the sensing path between transmit AP $t$ and receive AP $r$. We adopt the Swerling-I model, meaning that the RCS assumes only one value during the transmission and $\xi_{t,r}\sim\mathcal{CN}(0,\delta_{t,r}^{2})$. We assume that $\{\delta_{t,r}^{2}\}_{\forall t,r}$ are known and that RCSs belonging to different transmit-receiving pairs are statistically uncorrelated, that is
$\mathbb{E}\left\{\xi_{a,r}\xi_{a^{\prime},r^{\prime}}^{*}\right\}=\begin{cases}\delta_{a,r}^{2},&{\rm{if}}\leavevmode\nobreak\ a=a^{\prime},\leavevmode\nobreak\ r=r^{\prime},\\ 0,&\rm{otherwise}.\end{cases}$ (11)
In (10), $\mathbf{n}_{r}[\iota]$ is the receiver noise of AP $r$, assumed to have independent $\mathcal{CN}(0,\sigma^{2})$ entries. The vector $\mathbf{z}_{r}[\iota]$ represents the clutter as observed from AP $r$. The clutter vector is modeled as $\mathbf{z}_{r}[\iota]\sim\mathcal{CN}(\boldsymbol{0},\delta_{\mathbf{z}}^{2}\mathbf{R}_{r})$, where $\delta_{\mathbf{z}}^{2}$ is the power of the clutter cross section, which also includes the path loss effects, thus defining the CNR as $\textrm{CNR}=\delta_{\mathbf{z}}^{2}/\sigma^{2}$. The normalized clutter covariance matrix with trace being equal to $M$ is denoted by $\mathbf{R}_{r}$. This matrix is unknown, but its eigenspace is assumed to be known later in the detector design. For the sake of mathematical tractability, we assume that the clutter observations belonging to different APs and/or different timeslots are statistically uncorrelated. Under the LoS assumption for the channels involved in sensing, the target’s SCNR is defined as
@@
where $\overline{\mathbf{E}}_{a,r}=\overline{\mathbf{b}}_{r}\left(\overline{\mathbf{H}}_{a}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}\right)^{\mathsf{H}}$

### v3

@@ Abstract
In this paper, we investigate the performance of an integrated sensing and communication (ISAC) system within a cell-free massive multiple-input multiple-output (MIMO) system. Each access point (AP) operates in the millimeter-wave (mmWave) frequency band. The APs jointly serve the user equipments (UEs) in the downlink while simultaneously detecting a target through dedicated sensing beams directed toward a reconfigurable intelligent surface (RIS). Although the AP-RIS, RIS-target, and AP-target channels have both line-of-sight (LoS) and non-line-of-sight (NLoS) parts, only knowledge of the LoS paths is assumed to be available. A key contribution of this study is the consideration of clutter, which degrades the target detection performance if not handled. We propose an algorithm to alternatively optimize the transmit power allocation and the RIS phase-shift matrix, maximizing the target signal-to-clutter-plus-noise ratio (SCNR) while ensuring a minimum signal-to-interference-plus-noise ratio (SINR) for the UEs. Numerical results demonstrate that exploiting clutter subspace significantly enhances detection probability, particularly at high clutter-to-noise ratios, and reveal that an increased number of transmit side clusters impairs detection performance. Finally, we highlight the performance gains achieved using a dedicated sensing stream.
@@ I Introduction
Cell-free massive multiple-input multiple-output (MIMO) is a user-centric network infrastructure where a set of distributed access points serve multiple user equipments on the same time-frequency resources using joint processing techniques [1]. Over nearly a decade of research, it has been shown that cell-free massive MIMO improves both spectral and energy efficiency by providing macro-diversity and enhanced interference management compared to traditional cellular setups. As a result, it has become one of the key technology components considered for sixth-generation (6G) wireless networks [2]. Recently, cell-free massive MIMO architecture has also demonstrated potential benefits in integrated sensing and communications (ISAC) [3, 4]. ISAC allows for the efficient use of hardware and spectral resources through integration and coordination gains, compared to separate communication and sensing architectures [5]. Cell-free massive MIMO facilitates multi-static sensing through a central processing unit (CPU), to which the APs are connected via fronthaul links. Beyond cell-free massive MIMO and ISAC, another key 6G technology is reconfigurable intelligent surface (RIS)-aided communications. RISs have emerged as a promising technology in ISAC applications, enhancing the available spatial degrees of freedom. In [6], the sum rate of communication UEs is maximized under a worst-case sensing signal-to-noise ratio (SNR) constraint. An often overlooked aspect in ISAC research is the effect of clutter on target detection, which can severely degrade system performance if not properly addressed. Recent studies, such as [7], illustrate the impact of clutter in monostatic ISAC settings, underscoring the significance of clutter awareness at the receiving APs for reliable target detection.
@@
The contributions of this paper are outlined as follows
  • we investigate the ability of a RIS to improve target detection under transmitter-target obstruction. The need for precise channel state information (CSI) for target detection and the requirement for sensing-specific streams are explored within a millimeter-wave (mmWave) channel architecture.
  • A signal-to-clutter-and-noise ratio (SCNR)-maximizing alternating optimization (AO) algorithm is proposed, where the system assumes a line-of-sight (LoS) structure for the sensing links and optimizes the transmit power allocation and the RIS phase-shift matrix
  • numerical simulations reveal that clutter awareness and dedicated sensing streams significantly enhance the detection probability.
@@ II System Model
We consider the RIS-assisted cell-free massive MIMO ISAC network shown in Fig. 1. A set $\mathcal{T}$ of $T$ APs, each equipped with $M$ transmit antennas, serves $K$ single-antenna UEs in the downlink while sensing the potential presence of a target at a known position. Sensing is carried out over $\tau$ symbols belonging to the same channel coherence block. A second set $\mathcal{R}$ of $R$ APs collects the target echoes and performs target detection. The choice of geographically separated transmitting and receiving APs is motivated by the implementation challenges of full duplex operation, especially in terms of self-interference [8], and the interest in this setup for cell-free massive MIMO communications. We assume that the direct channel between the APs in $\mathcal{T}$ and the target is blocked and, thus, the formers reach the target through a RIS equipped with $N$ reflective elements, whose phase shifts are described by the vector $\boldsymbol{\theta}=[\theta_{1},\dots,\theta_{N}]^{\top}$, where $|\theta_{n}|=1$ for $n=1,\dots,N$. Each AP $t\in\mathcal{T}$ is connected to the RIS via the channel $\mathbf{H}_{t}\in\mathbb{C}^{N\times M}$, the reflected path connecting the RIS to the target is denoted by $\mathbf{c}\in\mathbb{C}^{N}$ and the UEs are connected to the APs in $\mathcal{T}$ via the channel $\{\mathbf{g}_{k,t}\}_{\forall k,t}\in\mathbb{C}^{M}$. The receiving APs in $\mathcal{R}$ are connected to the target via the channels $\{\mathbf{b}_{r}\}_{\forall r\in\mathcal{R}}\in\mathbb{C}^{M}$.
@@ II-A Channel models
Each of the previously mentioned channels is modelled in accordance with [9], which accurately captures the spatial sparsity that characterizes mmWave channels. Each channel consists of contributions from $C_{i}$ scattering clusters, i.e.,
@@
where the overline superscript indicates the LoS paths and the other terms correspond to the NLoS parts. The coefficients $\beta_{0,r}^{2}$, $\beta_{0}^{2}$, and $\beta_{0,t}^{2}$ are the corresponding LoS channel gains. We now remove the additional subscripts and superscripts except for $n$ as the following consideration applies regardless of the additional subscripts. Here, $\alpha_{n}$ is the complex channel gain of the $n$-th cluster. We assume each cluster scatters a sufficient number of rays such that $\alpha_{n}\sim\mathcal{CN}(0,\beta_{n}^{2})$ and the fading is mutually independent. The coefficients $\beta_{n}^{2}$ represent the path loss associated with the respective clusters. The geometry of the $n$-th cluster is described by $\boldsymbol{\omega}_{n}=[\psi_{n},\phi_{n}]$, where $\psi_{n}$ is the azimuth angle and $\phi_{n}$ is the elevation angle. Each AP is equipped with a uniform planar square array with a half-wavelength vertical and horizontal inter-element spacing such that the entries of $M$-dimensional steering vector are defined as $[\mathbf{a}_{M}(\boldsymbol{\omega}_{n})]_{m_{h}m_{v}}=e^{-j\pi(m_{h}\sin(\psi_{n})\cos(\phi_{n})+m_{v}\sin(\phi_{n}))}$,for $m_{h},m_{v}=0,\dots,\sqrt{M}-1$. We assume that perfect CSI regarding $\{\mathbf{g}_{k,t}\}_{\forall k,t}$ is available at the APs, as this is easily obtainable through standard estimation techniques. As for the sensing channels, the transmit APs know the position of the RIS and the target and, thus, the corresponding LoS channels.
@@ II-B Communication observation model
with $\mathbf{s}_{t}[\iota]=\mathbf{F}_{t}\mathbf{X}[\iota]\boldsymbol{\rho}$, where the matrix $\mathbf{X}[\iota]=\text{diag}(x_{1}[\iota],\ldots,x_{L}[\iota])$ is a diagonal matrix containing the downlink communication symbols and sensing symbol transmitted during timeslot $\iota$, where $\mathbb{E}\{|x_{k}[\iota]|^{2}\}=1$. The vector $\boldsymbol{\rho}=[\sqrt{\rho_{1}},\dots,\sqrt{\rho_{L}}]^{\top}$ contains the square root of the power allocated to each beam. Finally, $w_{k}[\iota]\sim\mathcal{CN}(0,\sigma_{k}^{2})$. In line with the centralized operation of cell-free massive MIMO, the communication precoding vectors are based on the transmit APs’ CSI [3] while the sensing precoder is based on the assumed LoS structure of the sensing channels and the known target location. To this end, we define $\mathbf{g}_{k}=[\mathbf{g}_{k,1}^{\top},\dots,\mathbf{g}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$ and $\mathbf{f}_{k}=[\mathbf{f}_{k,1}^{\top},\dots,\mathbf{f}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$. The communication precoding vectors for each UE are chosen using regularized zero forcing (RZF) as $\mathbf{f}_{k}=\overline{\mathbf{f}}_{k}/\left\lVert\overline{\mathbf{f}}_{k}\right\rVert$, where
$\overline{\mathbf{f}}_{k}=\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}$ (6)
where $\lambda$ is a regularization parameter. Let us define the matrix $\mathbf{G}=[\mathbf{g}_{1},\dots,\mathbf{g}_{K}]\in\mathbb{C}^{MT\times K}$ with all the UE channels. Then the sensing stream precoder is defined as $\mathbf{f}_{L,t}=\overline{\mathbf{f}}_{L,t}/\left\lVert\overline{\mathbf{f}}_{L,t}\right\rVert$
$\overline{\mathbf{f}}_{L,t}=\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}$ (7)
where $\overline{\mathbf{h}}=[\overline{\mathbf{h}}_{1}^{\top},\dots,\overline{\mathbf{h}}_{T}^{\top}]^{\top}$ and $\overline{\mathbf{h}}_{t}=\overline{\mathbf{H}}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}$ with $\boldsymbol{\Theta}=\text{diag}(\boldsymbol{\theta})$. We have projected $\overline{\mathbf{h}}$ onto the null space of the communication UEs’ channels to nullify the interference caused to the UEs. The communication performance is represented by the UEs’ spectral efficiency (SE), defined as ${\rm{SE}}_{k}=\log_{2}(1+{\rm{SINR}}_{k})$: under the assumption of perfect CSI on the UEs channels, the signal-to-interference-plus-noise ratio (SINR) is defined as
@@ II-C Sensing observation model
$\mathbf{y}_{r}[\iota]=\sum_{t\in\mathcal{T}}\xi_{t,r}\mathbf{E}_{t,r}\mathbf{s}_{t}[\iota]+\mathbf{z}_{r}[\iota]+\mathbf{n}_{r}[\iota],\vspace{-2mm}$ (9)
where $\xi_{t,r}$ is the radar cross section (RCS) associated with the sensing path between transmit AP $t$ and receive AP $r$. We adopt the Swerling-I model, meaning that the RCS assumes only one value during the transmission and $\xi_{t,r}\sim\mathcal{CN}(0,\delta_{t,r}^{2})$. We assume that $\{\delta_{t,r}^{2}\}_{\forall t,r}$ are known and that RCSs belonging to different transmit-receiving pairs are statistically uncorrelated, that is $\mathbb{E}\left\{\xi_{a,r}\xi_{a^{\prime},r^{\prime}}^{*}\right\}\neq 0\Leftrightarrow a=a^{\prime},\leavevmode\nobreak\ r=r^{\prime}$. In (9), $\mathbf{n}_{r}[\iota]$ is the receiver noise of AP $r$, assumed to have independent $\mathcal{CN}(0,\sigma^{2})$ entries. The vector $\mathbf{z}_{r}[\iota]$ represents the clutter as observed from AP $r$. The clutter vector is modelled as $\mathbf{z}_{r}[\iota]\sim\mathcal{CN}(\boldsymbol{0},\delta_{\mathbf{z}}^{2}\mathbf{R}_{r})$, where $\delta_{\mathbf{z}}^{2}$ is the power of the clutter cross-section, which also includes the path loss effects, thus defining the clutter-to-noise ratio (CNR) as $\textrm{CNR}=\delta_{\mathbf{z}}^{2}/\sigma^{2}$. The normalized clutter covariance matrix, with a trace equal to $M$, is denoted by $\mathbf{R}_{r}$. This matrix is unknown, but its eigenspace is assumed to be known later in the detector design. For the sake of mathematical tractability, we assume that the clutter observations belonging to different APs and/or different timeslots are statistically uncorrelated. Under the Los assumption for the channels involved in sensing, the target’s SCNR is defined as
@@
where $\overline{\mathbf{E}}_{a,r}=\overline{\mathbf{b}}_{r}\left(\overline{\mathbf{H}}_{a}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}\right)^{\mathsf{H}}$.